# Assessing Universal Relations for Rapidly Rotating Neutron Stars: Insights from an Interpretable Deep Learning Perspective

### v1

@@ Acknowledgements
G. P. wants to thank George Pappas for useful discussions. G. P. acknowledges that the research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship No. 20450). G. V. acknowledges partial funding support from HFRI Project No. 15940. The EoSs used in this work come from the ComPOSE database. We thank the team members of CompOSE for their development. The collection of neutron star data was conducted at the Department of Computer Science and Engineering, University of Ioannina, Greece. Especially, we would like to thank Professor Christophoros Nikou for generously providing access to the computing resources of the Impala cluster. The training of the proposed deep learning models was carried out using the computational resources of the Department of Physics at Aristotle University of Thessaloniki (AUTh), Greece. We gratefully acknowledge the support of the AUTh IT Center, as the results presented in this work were obtained using the university’s High-Performance Computing infrastructure.
@@
All necessary information supporting the findings of this study is included within the article. Representative code and notebook examples, along with the associated optimal parameters $\theta^{\star}$ (i.e., the trained model weights), are publicly available at: https://github.com/gregoryPapi/deep-universal-relations.git.
@@ Appendix A Equation of state models
Hybrid EoSs: DS(CMF)-1 Hybr, DS(CMF)-2 Hybr, DS(CMF)-3 Hybr, DS(CMF)-4 Hybr, DS(CMF)-5 Hybr, DS(CMF)-6 Hybr, DS(CMF)-7 Hybr, DS(CMF)-8 Hybr, DD2-FRG (2) flavors, DD2-FRG vector interaction-(2) flavors, QHC18, QHC19-B, QHC19-C, QHC19-D, QHC21T $A_{T}$, QHC21T $B_{T}$, QHC21T $C_{T}$, QHC21T $D_{T}$, QHC21 $A_{\chi}$, QHC21 $B_{\chi}$, QHC21 $C_{\chi}$, VQCD(APR) soft, VQCD(APR) intermidiate.
More details about EoS samples for each category, along with their specific theoretical descriptions and the associated microphysics compositions, can be found in Appendix B of Ref. [153] and the CompOSE database [160, 35].

### v2

@@ Acknowledgements
G. P. wants to thank George Pappas for useful discussions. G. P. acknowledges that the research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship No. 20450). G. V. acknowledges partial funding support from HFRI Project No. 15940. The EoSs used in this work come from the ComPOSE database. We thank the team members of CompOSE for their development. The collection of neutron star data was conducted at the Department of Computer Science and Engineering, University of Ioannina, Greece. Especially, we would like to thank Professor Christophoros Nikou for generously providing access to the computing resources of the Impala cluster. The training of the proposed deep learning models was carried out using the computational resources of the Department of Physics at Aristotle University of Thessaloniki (AUTh), Greece. We gratefully acknowledge the support of the AUTh IT Center, as the results presented in this work were obtained using the university’s High-Performance Computing infrastructure. Finally, we are grateful to the anonymous referee for their insightful comments and suggestions that improved the final version of this manuscript.
@@
All necessary information supporting the findings of this study is included within the article. Representative code and notebook examples, along with the associated optimal parameters $\theta^{\star}$ (i.e., the trained model weights), are publicly available in Ref. [193].
@@ Appendix A Equation of state models
Hybrid EoS models: DS(CMF)-1 Hybr, DS(CMF)-2 Hybr, DS(CMF)-3 Hybr, DS(CMF)-4 Hybr, DS(CMF)-5 Hybr, DS(CMF)-6 Hybr, DS(CMF)-7 Hybr, DS(CMF)-8 Hybr, DD2-FRG (2) flavors, DD2-FRG vector interaction-(2) flavors, QHC18, QHC19-B, QHC19-C, QHC19-D, QHC21T $A_{T}$, QHC21T $B_{T}$, QHC21T $C_{T}$, QHC21T $D_{T}$, QHC21 $A_{\chi}$, QHC21 $B_{\chi}$, QHC21 $C_{\chi}$, VQCD(APR) soft, VQCD(APR) intermidiate.
More details about EoS samples for each category, along with their specific theoretical descriptions and the associated microphysics compositions, can be found in Appendix B of Ref. [153] and the CompOSE database [160, 35]. Lastly, in Fig. 31, an EOS-color map is shown for the various parameter-space distributions demonstrated in Fig 2 of Sec. IV.
Figure 31: EoS-Color map used for the various parameter space representations given in Fig 2 of Sec. IV.